OpenStudy (anonymous):

help please!! 86 is 45% of what number?

OpenStudy (anonymous):

191

OpenStudy (accessdenied):

The percent just means, "how much of a whole do I have?" We can have 50%, or one-half, of a whole thing; or maybe 22%, which is 22/100 of a whole thing. In general, if a number is some percentage of another, we can say that (The whole number) * (The percentage) = (the number that is the percentage of the previous number) We can convert the percentage to a fraction which is more usable in math because x% = x/100. So, simple examples: 25 is 50% of 50, 2 is 20% of 10... 50 * 50% = 25 -> 50 * 1/2 = 25 10 * 20% = 2 -> 10 * 2/10 = 2 As in above, maybe 22% of something: What is 22% of 120? 120*22% = 120*22/100 = 26.4

OpenStudy (accessdenied):

For this problem... 86 is 45% of what number? So, some number, if we have 45% of it, we will end up with 86. If we call that number, say, #... # * 45% = 86. 45% = 45/100 45/100 * # = 86 Solve by multiplying both sides by 100/45 (to cancel with 45/100) # = 86 * 100/45 = 191.11... repeating